# HEST-1k: A Dataset for Spatial Transcriptomics and Histology Image Analysis

### v1

@@ 6 HEST for discovery
HEST-1k also enables the analysis of interactions and correlations between tissue morphology (as seen in H&E) and local gene expression (as provided in ST). Here, we showcase the capabilities of HEST-1k (1) by studying morphological correlates of expression changes in invasive breast cancer and (2) by visualizing tumor heterogeneity both on the morphological and molecular sides. Specifically, we focus on invasive ductal carcinoma (IDC) samples imaged with Xenium. Using CellViT nuclear segmentation and classification, we identified neoplastic nuclei (exemplified in two samples: Figure 2.a with n=168,033 nuclei and Appendix Figure 5.a with n=342,018 nuclei). We then overlay the WSI with the expression of specific genes, such as GATA3, a known prognostic gene in breast cancer Mehra et al. (2005)(Figure 2.b). This qualitatively shows that high GATA3 expression is associated with cancerous regions and reveals heterogeneity within invasive regions (e.g., the right-most region shows higher expression of GATA3 than the rest of the tumor, Figure 2.b). Using the nuclear segmentation, we can compute human-interpretable features related to nuclear size (area, perimeter, major axis length, minor axis length, and equivalent diameter), topology and shape (roundness, ellipticity, eccentricity, extent, and roughness), and cell distribution (cell density and crowdedness). A heatmap of the nuclear area of neoplastic cells also indicates morphological heterogeneity among neoplastic regions (Figure 2.c,d). Regions with a high nuclear area and elevated GATA3 expression notably overlap, suggesting that this tumor exhibits molecular heterogeneity, which to some degree is morphologically expressed.
@@
To investigate this hypothesis, we measured the Pearson correlation between the expression of GATA3 and nuclear area in neoplastic cells (Figure 2.e). We observe a significant correlation (R=0.47), which is also observed in other genes and morphological features (Figure 2.f). Overall, out of the 12 human-interpretable features we analyzed, we found the highest association with gene expression for size-related features, while features involving topology, shape, and cell distribution had a lower correlation (R<0.2). A similar analysis in another IDC sample (Appendix Figure 5.b,c) confirmed these observations. In particular, we found the highest associations between nuclear size and expression for the genes FLNB (R=0.49) and TPD52 (R=0.49), both involved in breast tumor growth and proliferation Bandaru et al. (2014); Ren et al. (2021), and FOXA1 (R=0.47), a known prognostic factor associated with better survival Badve et al. (2007); Wolf et al. (2007).
@@
Such analysis highlights how HEST-1k can be used for identifying fine-grained morphological correlates of expression. Additional analyses of morphological and molecular tumor heterogeneity can be leveraged at larger scale using similar approaches.
@@ 7 HEST for multimodal representation learning
Access to spatially-resolved expression–morphology pairs unlocks new directions for multimodal representation learning. Several problem statements can be explored, such as cross-modal alignment and retrieval, multimodal fusion, etc. Here, we fine-tune CONCH (strong performer on HEST-Benchmark) on five Xenium invasive breast cancer cases (four ductal and one lobular case) using contrastive alignment. We hypothesize that the resulting breast cancer-specific patch encoder, termed CONCH-FT, can better encode the underlying molecular landscape associated with disease-specific morphologies. To validate the hypothesis, CONCH-FT is benchmarked on an independent breast cancer cohort for molecular subtyping against its non-finetuned version.
@@
We evaluate the resulting CONCH-FT model to predict ER, PR, and HER2 expression status (binary) from WSIs in the BCNB dataset Xu et al. (2021) (n=1,058 WSIs). To generate a slide representation for a WSI, we take the average of the patch embedding in the WSI, which is subsequently mapped to the expression status using logistic regression (Table 2). We observe that CONCH-FT outperforms CONCH on most metrics, demonstrating that pan-tissue histology patch encoders can be further fine-tuned to obtain better tissue-specific patch encoders. This is further validated by the larger rank induced by the patch embedding space Garrido et al. (2023) for CONCH-FT, suggesting better expressivity of the patch embeddings. While these results are based on only five WSIs, we anticipate additional benefits when training with larger disease-specific cohorts.
Table 2: CONCH fine-tuning on invasive breast cancer. Logistic regression evaluation for ER/PR/HER2 expression status on BCNB (n=1,058 WSIs). A WSI is represented by the average of the patch embeddings within each WSI. We report the mean ± standard deviation computed over all folds (or patients) for ROC-AUC (AUC) and balanced accuracy (Bal.acc.). Best is bold.
@@ 8 Discussion
Summary. We assembled HEST-1k, a dataset comprising paired spatial transcriptomics, H&E-stained whole-slide images, and comprehensive metadata built from public and internal cohorts. HEST-1k includes 1,108 samples, encompassing 1.5 million spots and over 60 million cells. The scale and comprehensiveness of HEST-1k, supported by the HEST-Library, enable exploring directions such as biomarker discovery and multimodal representation learning. Additionally, motivated by the need for new diverse and challenging patch-level benchmarks, we curated the HEST-Benchmark, a set of ten tasks covering nine cancer types and ten organs for gene expression prediction from histology. HEST-Benchmark revealed new insights into the predictive capabilities of foundation models for histology, such as the added benefits of vision-language alignment Lu et al. (2024).
Limitations. Our study includes limitations. First, research data, such as those generated in spatial transcriptomic, are inherently noisy. While we tried to minimize “label” noise (e.g., by re-estimating image magnification and alignment, and unifying cancer samples using oncotree code taxonomy), staining and compression artifacts, varying acquisition protocols, among others can negatively impact the quality of HEST-1k. Second, batch effects (on both the imaging and transcriptomic sides) can be significant across samples, datasets, and technologies. This study does not quantify these batch effects or explore methods to mitigate them. Lastly, although the HEST-Library was designed for versatility, it cannot cover all existing formats and should rather be viewed as a blueprint for processing ST data in a consistent and unified manner.
Future work. Spatial transcriptomics is rapidly evolving with new datasets frequently published. As they become available, we will keep updating HEST-1k with new resources. This study merely starts to uncover the potential of HEST-1k for advancing translational research and biomarker discovery, and we plan to explore these capabilities further. Additionally, the prospects for multimodal representation learning with HEST-1k are promising and are expected to grow with the addition of more data.
@@ Acknowledgements
We thank Dr. Maxime Meylan for his insights and guidance on accessing data published in Meylan et al. (2022). S.J.W. was supported by the Helmholtz Association under the joint research school “Munich School for Data Science - MUDS” and the Add-on Fellowship of the Joachim Herz Foundation.
@@
    Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] Each claim: HEST-1k, HEST-Library, HEST-Benchmark, HEST for discovery and multimodal fine-tuning are supported by dedicated sections in the main text, in addition to supplementary information provided in the appendix. In addition, all the code to reproduce results is made available.
@@
    Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] In the abstract, we provide a link to access the HEST website and the HEST-Library on GitHub. HEST-Library includes a link to download all data to run the HEST-Benchmark. Finally, we provide all metadata associated with HEST-1k in a CSV as part of the supplementary material. Processed ST and WSIs will be made available upon acceptance.
@@
    Did you mention the license of the assets? [Yes] Metadata associated with HEST-1k includes the license under which data were originally published. We ensured that the reported license allowed for distributing and creating derivatives of the data.
@@ Appendix A Ethical considerations, intended usage and license
All resources provided as part of this study are strictly for research purposes and must not be utilized to support any diagnostic procedures. Users are hereby notified that the nuclear segmentation and classification components are derived from a publicly available model. Consequently, this model should not be regarded as the definitive standard, and users should exercise particular caution when utilizing this part of the dataset. Despite our efforts to exclude sensitive information, such as patient names, addresses, and social security numbers, users are expressly prohibited from attempting to reverse engineer the data to extract any confidential patient information. On the presumption that users will adhere to the aforementioned restrictions, we have not identified any potential adverse societal impacts that could arise from using HEST-1k.
The dataset is hosted on HuggingFace Dataset webpage. All instructions are provided in the main README of HEST-Library. From there, users can choose to download HEST-1k in its entirety, or a subset (e.g., only breast cancer samples). HEST-1k, HEST-Benchmark, and HEST library are released under the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 Deed) [^9] https://creativecommons.org/licenses/by-nc-sa/4.0/.
@@ Appendix B Background
This study connects two fields: (1) computational pathology, which primarily uses routinely acquired clinical data to determine outcomes such as disease diagnosis from H&E-stained digitized tissue sections, and (2) spatial transcriptomics, which has, so far, been confined to biological research and aims to identify new biomarkers predictive of disease progression or treatment response, among others.
@@ B.1 Computational pathology
Such frameworks have been shown to achieve better or similar performance than humans for Gleason grading in prostate cancerBulten et al. (2022), metastasis detection in lymph nodesBejnordi et al. (2017), determining the origin of a cancer of unknown primaryLu et al. (2021a), predicting heart transplant rejectionLipkova et al. (2022), among others.
@@ B.2 Spatial Transcriptomics (ST)
ST enables the measurement of gene activity and the mapping to its corresponding location on the tissue. In this study, we gather samples from two ST paradigms: sequencing-based (ST, Visium, Visium HD) and imaging-based (Xenium).
@@
Xenium: Xenium is an imaging-based spatial-profiling technology offering higher resolution compared to its sequencing counterpart, Visium. Xenium imaging-based paradigm breaks free from the resolution limitations of its sequencing-based counterpart, it directly performs in situ RNA capturing on tissue sections by imaging fluorescent RNA markers derived from padlock probes and rolling circle amplification chemistry, hence offering sub-cellular expression resolution. As of 2024, Xenium cannot perform whole transcriptome measurement and is limited to gene panels of up to 580 genes.
@@ C.1 HEST-1k
Table A1: HEST-1k data overview. All included samples include a license that allows sharing and redistributing. National Center for Biotechnology Information: NCBI.
@@
Normal Human Prostate (FFPE) Prostate Visium 1 17943
@@
10X Visium Spatial transcriptomics of murine colon at d14 (mucosa healing) in B cell sufficient-deficient mice Parigi et al. (2022) Bowel Visium 2 31,053
10X Visium Spatial transcriptomics of murine colon in steady state and during recovery after DSS colitis Parigi et al. (2022) Bowel Visium 2 31,053
A Spatial Transcriptomic atlas of the human kidney papilla identifies significant immune injury in patients with stone disease Canela et al. (2023) Kidney Visium 7 36,601
A cellular hierarchy in melanoma uncouples growth and metastasis Karras et al. (2022) Skin Visium 3 31,053
A new epithelial cell subpopulation predicts response to surgery, chemotherapy, and immunotherapy in bladder cancer Abdel-Hafiz et al. (2023); Gouin et al. (2021) Bladder Visium 4 33,538
A novel model of binge ethanol exposure reveals enhanced neurodegeneration with advanced age Anton et al. (2024) Brain Visium 4 32,285
A single-cell transcriptomic analysis of endometriosis Fonseca et al. (2023) Uterus Visium 2 36,601
Distinct mesenchymal cell states mediate prostate cancer progression Pakula et al. (2024) Prostate Visium 2 32,589
Epithelial Plasticity and Innate Immune Activation Promote Lung Tissue Remodeling following Respiratory Viral Infection Beppu et al. (2023) Lung Visium 1 32,285
Gene expression within a human choroidal neovascular membrane using spatial transcriptomics Voigt et al. (2023) Eye Visium 5 36,601
Genome-wide Spatial Expression Profiling in Formalin-fixed Tissues Gracia Villacampa et al. (2021) Kidney Visium 14 33,538
High resolution mapping of the tumor microenvironment using integrated single-cell, spatial and in situ analysis Janesick et al. (2023) Breast Xenium 4 541
Identification of TREM1+CD163+ myeloid cells as a deleterious immune subset in HCC Giraud et al. (2022) Liver Visium 2 36,601
Integration of spatial and single cell transcriptomics localizes epithelial-immune cross-talk in kidney injury Ferreira et al. (2021) Kidney Visium 4 33,538
Molecular Atlas of the Adult Mouse Brain Ortiz et al. (2020) Brain Spatial Transcriptomics 75 23,371
@@
SARS-CoV-2 Niches in Human Placenta Revealed by Spatial Transcriptomics Barrozo et al. (2023) Uterus Visium 16 36,612
Schwann Cells Shape Tumor Cells and Cancer-Associated Fibroblasts in the Pancreatic Ductal Adenocarcinoma Microenvironment Xue et al. (2023) Pancreas Visium 4 20,615
Single Cell and Spatial Analysis of Human Squamous Cell Carcinoma Ji et al. (2020) Skin Spatial Transcriptomics 12 17,138
Single cell profiling of primary and paired metastatic lymph node tumors in breast cancer patients Liu et al. (2022) Lymph node Visium 4 33,931
@@
Single-cell and spatial transcriptomics characterisation of the immunological landscape in the healthy and PSC human liver Andrews et al. (2024) Liver Visium 4 36,601
Single-nucleus Ribonucleic Acid-sequencing and Spatial Transcriptomics Reveal the Cardioprotection of Shexiang Baoxin Pill (MUSKARDIA) in Mice with Myocardial Ischemia-Reperfusion Injury Lin et al. (2023) Heart Visium 2 32,285
Spatial Multimodal Analysis: MALDI-MSI and Spatial Transcriptomics within the same tissue section Vicari et al. (2023) Brain Visium 19 32,285
Spatial RNA sequencing of regenerating mouse hindlimb muscle McKellar et al. (2021) Muscle Visium 3 33,217
Spatial Total RNA-Sequencing of regenerating mouse hindlimb muscle and Type 1-Lang reovirus-infected mouse heart McKellar et al. (2023) Muscle Visium 7 55,414
Spatial localization with Spatial Transcriptomics for an atlas of healthy and injured cell states and niches in the human kidney Lake et al. (2023) Kidney Visium 23 33,538
Spatial sequencing of Foreign body granuloma Krausgruber et al. (2023) None Visium 1 15,524
Spatial transcriptomics landscape of non-communicable inflammatory skin diseases Schäbitz et al. (2022) Skin Visium 59 20,613
Spatial transcriptomics of adenoid cystic carcinoma of the lacrimal gland Moeyersoms et al. (2023) Eye Visium 1 17,943
@@
Spatial transcriptomics reveal unnresolved wound repair as potential driver of PFA Ependymoma progression Fu et al. (2023) Brain Visium 14 36,601
Spatiotemporal dynamics of molecular pathology in amyotrophic lateral sclerosis Silas Maniatis et al. (2019) Spinal cord Spatial Transcriptomics 302 12,572
The neurons that restore walking after paralysis Kathe et al. (2022) Spinal cord Visium 16 22,127
Visium spatial transcriptomics analysis of lacrimal gland during chronic inflammation progression Mauduit et al. (2022) Eye Visium 4 32,285
YAP Drives Assembly of a Spatially Colocalized Cellular Triad Required for Heart Renewal Li et al. (2024) Heart Visium 2 32,285
Zika virus co-opts miRNA networks to persist in placental microenvironments detected by spatial transcriptomics Barrozo et al. (2024) Placenta Visium 8 32,298
Mouse model Heptablastoma spatial transcriptomics Pilet et al. (2023) Liver Visium 10 31,053
@@
A spatiotemporal organ-wide gene expression and cell atlas of the developing human heart Asp et al. (2019) Heart Spatial Transcriptomics 19 39,739
Integrating spatial gene expression and breast tumour morphology via deep learning He et al. (2020) Breast Spatial Transcriptomics 68 16,744
Spatial deconvolution of HER2-positive breast cancer delineates tumor-associated cell type interactions Andersson et al. (2021) Breast Spatial Transcriptomics 36 15,045
Visualization and analysis of gene expression in tissue sections by spatial transcriptomics Ståhl et al. (2016a) Brain Spatial Transcriptomics 16 16,573

### v2

@@ 6 HEST for biomarker exploration
HEST-1k also enables the analysis of interactions and correlations between tissue morphology (as seen in H&E) and local gene expression (as provided in ST). Here, we showcase the capabilities of HEST-1k (1) by studying morphological correlates of expression changes in invasive breast cancer and (2) by visualizing tumor heterogeneity both on the morphological and molecular sides. Specifically, we focus on invasive ductal carcinoma (IDC) samples imaged with Xenium. Using CellViT nuclear segmentation and classification, we identified neoplastic nuclei (exemplified in two samples: Figure 3.a with n=168,033 nuclei and Appendix Figure 6.a with n=342,018 nuclei). We then overlay the WSI with the expression of specific genes, such as GATA3, a known prognostic gene in breast cancer [mehra2005identification](Figure 3.b). This qualitatively shows that high GATA3 expression is associated with cancerous regions and reveals heterogeneity within invasive regions (e.g., the right-most region shows higher expression of GATA3 than the rest of the tumor, Figure 3.b). Using the nuclear segmentation, we can compute human-interpretable features related to nuclear size (area, perimeter, major axis length, minor axis length, and equivalent diameter), topology and shape (roundness, ellipticity, eccentricity, extent, and roughness), and cell distribution (cell density and crowdedness). A heatmap of the nuclear area of neoplastic cells also indicates morphological heterogeneity among neoplastic regions (Figure 3.c,d). Regions with a high nuclear area and elevated GATA3 expression notably overlap, suggesting that this tumor exhibits molecular heterogeneity, which to some degree is morphologically expressed.
@@
To investigate this hypothesis, we measured the Pearson correlation between the expression of GATA3 and nuclear area in neoplastic cells (Figure 3.e). We observe a moderate correlation (R=0.47, P-value$<10^{-4}$), which is also observed in other genes and morphological features (Figure 3.f). Overall, out of the 12 human-interpretable features we analyzed, we found the highest association with gene expression for size-related features, while features involving topology, shape, and cell distribution had a lower correlation (R<0.2). A similar analysis in another IDC sample (Appendix Figure 6.b,c) further asserted these observations. In particular, we found the highest associations between nuclear size and expression for the genes FLNB (R=0.45, P-value$<10^{-4}$) and TPD52 (R=0.47, P-value$<10^{-4}$), both involved in breast tumor growth and proliferation [bandaru2014targeting, ren2021tumor], and FOXA1 (R=0.47, P-value$<10^{-4}$), a known prognostic factor associated with better survival [badve2007foxa1, wolf2007foxa1].
@@
Such analysis highlights how HEST-1k can be used to identify fine-grained morphological correlates of expression. Similar approaches can be used to characterize morphological and molecular tumor heterogeneity at a larger scale.
Figure 3: HEST for biomarker exploration: Analysis of an invasive ductal carcinoma sample imaged with Xenium. a. IDC Xenium sample with neoplastic nuclei overlaid in red ($n_{c}=168,033$ detected nuclei). Gray scale bar represents 2 mm. b. Heatmap of Xenium expression of gene GATA3. Blue and red values indicate above and below the mean (in white), respectively. c. Heatmap of neoplastic nuclear area. d. Four randomly selected regions with CellViT segmentation of the neoplastic nuclei. Black scale bar represents 30 $\mu$m. e., f. Correlation between nuclear area and GATA3, and minor axis length and MYBPC1.
@@ 7 HEST for multimodal representation learning
Access to spatially-resolved expression–morphology pairs unlocks new directions for multimodal representation learning. Several problem statements can be explored, such as cross-modal alignment and retrieval, multimodal fusion, etc. Here, we fine-tune CONCH [lu2024towards] (ViT-Base model) on five Xenium invasive breast cancer cases (four ductal and one lobular case) using multimodal contrastive alignment. We hypothesize that the resulting breast cancer-specific patch encoder, termed CONCH-FT, can better encode the underlying molecular landscape associated with disease-specific morphologies. To validate the hypothesis, CONCH-FT is benchmarked on an independent breast cancer cohort for molecular subtyping against its non-finetuned version.
@@
We evaluate the resulting CONCH-FT model to predict ER, PR, and HER2 expression status (binary) from WSIs in the BCNB dataset [xu2021predicting] (n=1,058 WSIs). To generate a slide representation for a WSI, we take the average of the patch embedding in the WSI (mean pooling), which is subsequently mapped to the expression status using logistic regression (Table 2). The simple mean pooling approach to embedding the slide without additional fine-tuning on the downstream tasks highlights the expressivity of the learned latent space. We observe that CONCH-FT outperforms CONCH on most metrics, demonstrating that pan-tissue histology patch encoders can be further fine-tuned to obtain better tissue-specific patch encoders. This is further validated by the larger rank induced by the patch embedding space [garrido2023rankme] for CONCH-FT, suggesting better expressivity of the patch embeddings. While these results are based on only five paired WSIs, we anticipate additional benefits when training with larger disease-specific cohorts.
Table 2: CONCH fine-tuning on invasive breast cancer. Logistic regression evaluation for ER/PR/HER2 status on BCNB (binary task, n=1,058 WSIs). A WSI is represented by the average of the patch embeddings within each WSI. We report the mean ± standard deviation computed over all folds (or patients) for ROC-AUC (AUC) and balanced accuracy (Bal.acc.). Best is bold.
@@ 8 Discussion
Summary. We assembled HEST-1k, a dataset comprising paired spatial transcriptomics, H&E-stained whole-slide images, and comprehensive metadata built from public and internal cohorts. HEST-1k includes 1,229 samples, encompassing 2.1 million spots and over 76 million cells. The scale and comprehensiveness of HEST-1k, supported by the HEST-Library, enable exploring directions such as biomarker exploration and multimodal representation learning. Additionally, motivated by the need for new, diverse, and challenging patch-level benchmarks, we curated the HEST-Benchmark, a set of nine tasks covering eight cancer types and nine organs for gene expression prediction from histology. The HEST-Benchmark revealed data and model scaling laws across 11 foundation models of different dimensions and pretraining scale [lu2024towards].
Limitations. Our study includes a few limitations. First, research data, such as those generated in spatial transcriptomic, are inherently noisy. While we tried to minimize “label” noise (e.g., by re-estimating image magnification and alignment, and unifying cancer samples using oncotree code taxonomy), staining and compression artifacts, varying acquisition protocols, among others, can negatively impact the quality of HEST-1k. Second, batch effects (on both the imaging and transcriptomic sides) can be significant across samples, datasets, and technologies. While this study does not explore batch effects quantification or mitigation, we provide a set of helpers in HEST-Library to let users explore this direction. Lastly, although the HEST-Library was designed for versatility, it cannot cover all existing formats and should rather be viewed as a blueprint for processing ST data in a consistent and unified manner.
Future work. Spatial transcriptomics is rapidly evolving, with new datasets frequently published. As they become available, we will keep updating HEST-1k with new resources. This study merely starts to uncover the potential of HEST-1k for advancing translational research and biomarker exploration, and we plan to explore these capabilities further. Additionally, the prospects for multimodal representation learning with HEST-1k are promising and are expected to grow with the addition of more data.
@@ Acknowledgements
We thank Dr. Maxime Meylan for his insights and guidance on accessing data published in [meylan2022tertiary]. We thank Rushin Gindra for his support in inspecting HEST-1k data, reporting issues, and providing references. HEST is supported by the Brigham and Women’s Hospital (BWH) President’s Fund, Mass General Hospital (MGH) Pathology, and the National Institute of Health (NIH) National Institute of General Medical Sciences (NIGMS) through R35GM138216. S.J.W. is supported by the Helmholtz Association under the joint research school “Munich School for Data Science - MUDS” and the Add-on Fellowship of the Joachim Herz Foundation.
@@
    Do the main claims made in the abstract, and introduction accurately reflect the paper’s contributions and scope? [Yes] Each claim: HEST-1k, HEST-Library, HEST-Benchmark, HEST for biomarker exploration, and multimodal fine-tuning are supported by dedicated sections in the main text, in addition to supplementary information provided in the appendix. In addition, all the code to reproduce results is made available.
@@
    Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] In the abstract, we provide a link to access the HEST page on GitHub. HEST-Library includes a link to download all data and to run the HEST-Benchmark. Finally, we provide all metadata associated with HEST-1k in a CSV as part of the supplementary material.
@@
    Did you mention the license of the assets? [Yes] Metadata associated with HEST-1k includes the license under which data were originally published. We ensured that the reported license allowed the distribution and creation of derivatives of the data.
@@ Appendix A Ethical considerations, intended usage and license
All resources provided as part of this study are strictly for research purposes and must not be utilized to support any diagnostic procedures. Users are hereby notified that the nuclear segmentation and classification components are derived from a publicly available model. Consequently, this model should not be regarded as the definitive standard, and users should exercise particular caution when utilizing this part of the dataset. Despite our efforts to exclude sensitive information, such as patient names, addresses, and social security numbers, users are expressly prohibited from attempting to reverse engineer the data to extract confidential patient information. In the presumption that users will adhere to the aforementioned restrictions, we have not identified any potential adverse social impacts that could arise from the use of HEST-1k.
The dataset is hosted on the HuggingFace Dataset webpage. All instructions are provided in the main README of HEST-Library. From there, users can choose to download HEST-1k in its entirety or a subset (e.g., only breast cancer samples). The HEST-1k, HEST-Benchmark, and HEST library are released under the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 Deed) [^11] https://creativecommons.org/licenses/by-nc-sa/4.0/.
@@ Appendix B Background
This study connects two fields: (1) computational pathology, which primarily uses routinely acquired clinical data to determine outcomes such as disease diagnosis from H&E-stained digitized tissue sections, and (2) spatial transcriptomics, which so far has been confined to biological research and aims to identify new biomarkers predictive of disease progression or response to treatment, among others.
@@ B.1 Computational pathology
Such frameworks have been shown to achieve better or similar performance than humans for Gleason grading in prostate cancer[bulten2022artificial], metastasis detection in lymph nodes[bejnordi2017diagnostic], determining the origin of a cancer of unknown primary[lu2021ai], predicting heart transplant rejection[lipkova2022deep], among others.
@@ B.2 Spatial transcriptomics (ST)
ST enables the measurement of gene activity and the mapping of its corresponding location in the tissue. In this study, we collected samples from two ST paradigms: sequencing-based (ST, Visium, Visium HD) and imaging-based (Xenium).
@@
Xenium: Xenium is an imaging-based spatial profiling technology that offers in situ RNA capturing on tissue sections by imaging fluorescent RNA markers derived from padlock probes and rolling circle amplification chemistry. This approach provides the exact 2D location of each measured transcript. As of 2024, Xenium cannot perform whole transcriptome measurements and is limited to gene panels of up to 5,000 genes.
@@ C.1 HEST-1k
Table A1: HEST-1k data overview. All samples include a license that allows sharing and redistributing. National Center for Biotechnology Information: NCBI.
@@
Characterization of immune cell populations in the tumor microenvironment of colorectal cancer using high definition spatial profiling [oliveira2024characterization] Bowel Mixed 8 18,085
@@
FFPE Human Prostate Adenocarcinoma with 5K Human Pan Tissue and Pathways Panel Prostate Xenium 1 10,006
FFPE Human Skin Primary Dermal Melanoma with 5K Human Pan Tissue and Pathways Panel Skin Xenium 1 10,017
@@
Fresh Frozen Mouse Brain Hemisphere with 5K Mouse Pan Tissue and Pathways Panel Brain Xenium 1 13,780
@@
Normal Human Prostate (FFPE) Prostate Visium 1 17,943
@@
Preview Data: FFPE Human Lymph Node with 5K Pan Tissue and Pathways Panel Lymphoid Xenium 1 11,094
@@
Visium HD Spatial Gene Expression Library, Mouse Kidney (FFPE) Kidney Visium HD 1 19,059
Visium HD Spatial Gene Expression Library, Mouse Embryo (FFPE) Embryo Visium HD 1 19,059
Visium HD Spatial Gene Expression Library, Human Pancreas (FFPE) Pancreas Visium HD 1 18,085
@@
10X Visium Spatial transcriptomics of murine colon at d14 (mucosa healing) in B cell sufficient-deficient mice [parigi2022spatial] Bowel Visium 2 31,053
10X Visium Spatial transcriptomics of murine colon in steady state and during recovery after DSS colitis [parigi2022spatial] Bowel Visium 2 31,053
A Spatial Transcriptomic atlas of the human kidney papilla identifies significant immune injury in patients with stone disease [canela2023spatially] Kidney Visium 7 36,601
A cellular hierarchy in melanoma uncouples growth and metastasis [karras2022cellular] Skin Visium 3 31,053
A new epithelial cell subpopulation predicts response to surgery, chemotherapy, and immunotherapy in bladder cancer [abdelhafiz2023ychromosome, gouin2021ncadherin] Bladder Visium 4 33,538
A novel model of binge ethanol exposure reveals enhanced neurodegeneration with advanced age [anton2024binge] Brain Visium 4 32,285
A single-cell transcriptomic analysis of endometriosis [fonseca2023single] Uterus Visium 2 36,601
Distinct mesenchymal cell states mediate prostate cancer progression [pakula2024distinct] Prostate Visium 2 32,589
Epithelial Plasticity and Innate Immune Activation Promote Lung Tissue Remodeling following Respiratory Viral Infection [beppu2023epithelial] Lung Visium 1 32,285
Image-based spatial transcriptomics identifies molecular niche dysregulation associated with distal lung remodeling in pulmonary fibrosis [Vannan2023.12.15.571954] Lung Xenium 20 17,145
Gene expression within a human choroidal neovascular membrane using spatial transcriptomics [voigt2023gene] Eye Visium 5 36,601
Genome-wide Spatial Expression Profiling in Formalin-fixed Tissues [graciaVillacampa2021genome] Kidney Visium 14 33,538
High-resolution mapping of the tumor microenvironment using integrated single-cell, spatial, and in situ analysis [janesick2023high] Breast Xenium 4 541
Identification of TREM1+CD163+ myeloid cells as a deleterious immune subset in HCC [giraud2022trem] Liver Visium 2 36,601
Integration of spatial and single cell transcriptomics localizes epithelial-immune cross-talk in kidney injury [ferreira2021integration] Kidney Visium 4 33,538
Molecular Atlas of the Adult Mouse Brain [ortiz2020molecular] Brain Spatial Transcriptomics 75 23,371
@@
SARS-CoV-2 Niches in Human Placenta Revealed by Spatial Transcriptomics [barrozo2023sars] Uterus Visium 16 36,612
Schwann Cells Shape Tumor Cells and Cancer-Associated Fibroblasts in the Pancreatic Ductal Adenocarcinoma Microenvironment [xue2023schwann] Pancreas Visium 4 20,615
Single Cell and Spatial Analysis of Human Squamous Cell Carcinoma [ji2020multimodal] Skin Spatial Transcriptomics 12 17,138
Single-cell profiling of primary and paired metastatic lymph node tumors in breast cancer patients [liu2022single] Lymph node Visium 4 33,931
@@
Single-cell and spatial transcriptomics characterization of the immunological landscape in the healthy and PSC human liver [andrews2024single] Liver Visium 4 36,601
Single-nucleus Ribonucleic Acid-sequencing and Spatial Transcriptomics Reveal the Cardioprotection of Shexiang Baoxin Pill (MUSKARDIA) in Mice with Myocardial Ischemia-Reperfusion Injury [lin2023single] Heart Visium 2 32,285
Spatial Multimodal Analysis: MALDI-MSI and Spatial Transcriptomics within the same tissue section [vicari2023spatial] Brain Visium 19 32,285
Spatial RNA sequencing of regenerating mouse hindlimb muscle [mckellar2021large] Muscle Visium 3 33,217
Spatial Total RNA-Sequencing of regenerating mouse hindlimb muscle and Type 1-Lang reovirus-infected mouse heart [mcKellar2023spatial] Muscle Visium 7 55,414
Spatial localization with Spatial Transcriptomics for an atlas of healthy and injured cell states and niches in the human kidney [lake2023an] Kidney Visium 23 33,538
Spatial sequencing of Foreign body granuloma [krausgruber2023single] None Visium 1 15,524
Spatial transcriptomics landscape of non-communicable inflammatory skin diseases [schabitz2022spatial] Skin Visium 59 20,613
Spatial transcriptomics of adenoid cystic carcinoma of the lacrimal gland [moeyersoms2023spatial] Eye Visium 1 17,943
@@
Spatial transcriptomics reveal unnresolved wound repair as potential driver of PFA Ependymoma progression [fu2023spatial] Brain Visium 14 36,601
Spatiotemporal dynamics of molecular pathology in amyotrophic lateral sclerosis [maniatis2019spatiotemporal] Spinal cord Spatial Transcriptomics 302 12,572
The neurons that restore walking after paralysis [kathe2022neurons] Spinal cord Visium 16 22,127
Visium spatial transcriptomics analysis of lacrimal gland during chronic inflammation progression [mauduit2022spatial] Eye Visium 4 32,285
YAP Drives Assembly of a Spatially Colocalized Cellular Triad Required for Heart Renewal [li2024yap] Heart Visium 2 32,285
Zika virus co-opts miRNA networks to persist in placental microenvironments detected by spatial transcriptomics [barrozo2024zika] Placenta Visium 8 32,298
Mouse model Heptablastoma spatial transcriptomics [pilet2023preneoplastic] Liver Visium 10 31,053
@@
A spatiotemporal organ-wide gene expression and cell atlas of the developing human heart [asp2019spatiotemporal] Heart Spatial Transcriptomics 19 39,739
Integrating spatial gene expression and breast tumour morphology via deep learning [he2020integrating] Breast Spatial Transcriptomics 68 16,744
Spatial deconvolution of HER2-positive breast cancer delineates tumor-associated cell type interactions [andersson2021spatial] Breast Spatial Transcriptomics 36 15,045
Visualization and analysis of gene expression in tissue sections by spatial transcriptomics [stahl2016visualization] Brain Spatial Transcriptomics 16 16,573